OpenStudy (mven):

Find the volume of a cylinder of height 6" and base with circumference 16"

OpenStudy (mven):

V = pir2h right?

OpenStudy (igreen):

Yep

OpenStudy (igreen):

Base circumference is 16..so find the radius.

OpenStudy (igreen):

\(\sf C = 2 \pi r\) Plug in what we know: \(\sf 16 = 2 (3.14)r\) Solve for 'r', multiply 2 * 3.14 and divide it to 16.

OpenStudy (mven):

0.3925

OpenStudy (igreen):

Check again

OpenStudy (mven):

2.54?

OpenStudy (mven):

divide it by 16 or 16 by it?

OpenStudy (igreen):

2.55 if you want to round.

OpenStudy (mven):

ahh ok then V = 2.55(2)*6

OpenStudy (igreen):

Not quite. \(V = (3.14)(2.55^2)(6)\)

OpenStudy (mven):

ahh yes forgot pi

OpenStudy (igreen):

Does it want you to simplify pi?

OpenStudy (mven):

no, its multiple choice and it gives me 96pi 192pi 384pi

OpenStudy (igreen):

Oh, okay..so forget about pi. \(\sf V = (2.55^2)(6)\)

OpenStudy (mven):

39.015

OpenStudy (igreen):

Yep that's our answer.

OpenStudy (mven):

not one of the choices?

OpenStudy (igreen):

"Find the volume of a cylinder of height 6" and base with circumference 16"" Is that 'with circumference of 16 pi'?

OpenStudy (mven):

yeah sorry it didnt copy over

OpenStudy (igreen):

Okay, so we have to find the radius over again. \(\sf C = 2 \pi r\) Plug in what we know: \(\sf 16\pi = 2 \pi r\) pi cancels out: \(\sf 16 = 2r\) Divide 2 to both sides, what's 16 / 2?

OpenStudy (mven):

8

OpenStudy (igreen):

Yes, now plug that into the formula again. \(\sf V = \pi r^2 h\) \(\sf V = (8)^2(6)\) Simplify

OpenStudy (mven):

so 384?

OpenStudy (igreen):

Yep, you got it.

OpenStudy (mven):

awesome thanks you